# Onsager vortex clusters on a sphere

### v1

@@ Abstract
We study Onsager vortex clustered states in a shell-shaped superfluid containing a large number of quantum vortices. In the incompressible limit and at low temperatures, the relevant problem can be boiled down to the statistical mechanics of neutral point vortices confined on a sphere. We analyze rotation free vortex clustered states within the mean field theory in the microcanonical ensemble. We find that the sandwich state, which involves the separating of vortices with opposite circulation and the clustering of vortices with same circulation around the poles and the equator, is the maximum entropy vortex distribution, subject to zero angular momentum constraint. The dipole momentum vanishes for the sandwich state and the quadrupole tensor serves as an order parameter to characterize the vortex cluster structure. For given finite angular momentum, the equilibrium vortex distribution forms a dipole structure, i.e., vortices with opposite sign are separated and are accumulated around the south and north pole, respectively. The conditions for the onset of clustering, and the exponents associated with the quadrupole moment and the dipole moment as functions of energy, are obtained within the mean field theory. At large energies, we obtain asymptotically exact vortex density distributions using the stereographic projection method, which give rise the parameter bounds for the vortex clustered states. The analytical predictions are in excellent agreement with microcanonical Monte Carlo simulations.
@@ I Introduction
Previous studies on Onsager vortices mainly focus on systems confined on a flat region. Recent experimental realizations of BECs Aveline et al. (2020) and ultracold atomic bubbles Carollo et al. (2022) in International Space Station open a possibility to investigate a bubble trapped superfluid and vortex physics on curved surfaces experimentally. Curvature and topology of a surface have non-trivial effects on BECs Tononi and Salasnich (2019); Tononi et al. (2020); Tononi and Salasnich (2023), vortex dynamics Bogomolov (1978); CROWDY (2004); Hally (1980); Polvani and Dritschel (1993); Kimura (1999); Newton (2001); Alobaidi et al. (2006); Aref (2007); Vitelli and Turner (2004); Turner et al. (2010); Dritschel and Boatto (2015); Dritschel et al. (2015); Doshi and Gromov (2021); Bereta et al. (2021); Caracanhas et al. (2022); Padavić et al. (2020); Saito and Hayashi (2023); Li and Efimkin (2023); Xiong and Yu (2024), and the Berezinskii-Kosterlitz-Thouless (BKT) transition Tononi et al. (2022). The formation and structure of Onsager vortices also strongly depend on the surface geometry. Numerical simulations of the Gross-Pitaevskii equation show that, different from the giant dipole configuration for vortices confined to a discs Simula et al. (2014); Yu et al. (2016), in a spherical shell trap uniformly distributed neutral vortices form a quadruple cluster through the evaporative heating process Kanai and Guo (2021). However, it is not clear whether the quadruple clustered state is the statistically most favored state for given energy, angular momentum and vortex number.
@@
In this paper, we systematically investigate rotation-free clustered states of a neutral vortex system confined on a sphere. We extend the mean field theory of vortex statistical mechanics developed by Joyce and Montgomery Joyce and Montgomery (1973) to a closed surface. In particular, we analyze the self-consistent equation of the stream function on a sphere near the uniform state and in the high energy limit. We find that, under constraint of zero angular momentum, the sandwich state where vortices with opposite sign are distributed around the poles and the equator respectively, is the maximum entropy state and has higher statistical weight than the quarupole state. This theoretical prediction is conformed by Monte Carlo simulations.
II Point vortices on a closed surface
@@
For a superfluid on a plane, the circulation of a quanutm vortex is quantized in units of circulation quantum $\kappa\equiv 2\pi\hbar/m$ Pitaevskii and Stringari (2016), and the vorticity has a singularity at the vortex core $\mathbf{r}_{i}$: $\omega(\mathbf{r})=\nabla\times\mathbf{u}=\kappa_{i}\delta(\mathbf{r}-\mathbf{r}_{i})=\kappa\sigma_{i}\delta(\mathbf{r}-\mathbf{r}_{i})$ with sign $\sigma_{i}=\pm 1$ for singly charged vortices. Here $m$ is the atomic mass and $\mathbf{u}$ is the superfluid velocity. The quantization ensures that the vorticity of a quantum vortex concentrates around the core region Eyink and Sreenivasan (2006). Hence when the mean separation between quantum vortices $\ell$ is much larger than the vortex core size $\xi$ Fetter (1998); Barenghi and Parker (2016), the point vortex model (PVM) describes well the dynamics of quantum vortices in the incompressible limit Onsager (1949); Eyink and Sreenivasan (2006); Novikov (1975); Aref et al. (1999), provided vortex annihilation can be neglected. The PVM also describes 2D superfluid transition at finite temperature, which is the celebrated BKT transition Kosterlitz and Thouless (1973); Kosterlitz (1974).
@@
On curved surfaces, a covariant PVM has been well formulated Hally (1980); Kimura (1999); Dritschel and Boatto (2015). Here we summarize the main results of the PVM on a closed surface. In isothermal coordinates, the metric on the surface is
$ds^{2}=g_{\mu\nu}dx^{\mu}dx^{\nu}=h(x^{1},x^{2})[(dx^{1})^{2}+(dx^{2})^{2}],$ (1)
where $g_{12}=g_{21}=0$, $g_{11}=g_{22}=h(x^{1},x^{2})$, and $h(x^{1},x^{2})$ is a positive function and exists locally for 2D surfaces Chern (1955). For incompressible fluids, the component fluid velocity $u_{a}=\epsilon_{a}^{\,\,\,b}\nabla_{b}\psi$, where $\psi$ is the stream function, $\nabla_{b}$ is the Levi-Civita covariant derivative, $\epsilon^{\beta}_{\,\,\,\,\alpha}=g^{\beta\gamma}\epsilon_{\gamma\alpha}$, $\epsilon_{\gamma\alpha}=\sqrt{\det g_{\mu\nu}}\,\tilde{\epsilon}_{\gamma\alpha}$ is the Levi-Civita tensor and $\tilde{\epsilon}_{\gamma\alpha}$ is the Levi-Civita symbol. The vorticity $\omega=\epsilon^{\mu}_{\,\,\,\,\nu}\nabla_{\mu}u^{\nu}=-\Delta\psi$, where $\Delta\equiv\nabla^{\mu}\nabla_{\mu}=(1/\sqrt{g})\partial_{i}\left(\sqrt{g}g^{ij}\partial_{j}\right)=h^{-1}(x^{1},x^{2})(\partial^{2}_{x^{1}}+\partial^{2}_{x^{2}})$ is the Laplace-Bertrami operator and $g=\det g_{ij}$.
Let us also introduce complex coordinates $z=x^{1}+ix^{2}$, $\partial_{z}\equiv(\partial_{1}-i\partial_{2})/2$, $\partial_{\bar{z}}\equiv(\partial_{1}+i\partial_{2})/2$ and complex velocity $u\equiv u^{1}-iu^{2}$. In complex notations, $u=2ih^{-1}\partial_{z}\psi$ and the stream function
$\psi(z)=\kappa\sum_{i}\sigma_{i}G(z,z_{i}),$ (2)
where $z_{i}$ is the position of the $i$th vortex, $G(z,z_{i})$ is the Green’s function satisfying Dritschel and Boatto (2015)
$\Delta G(z,z_{i})=-\delta_{z,z_{i}}+1/\Omega,$ (3)
$\Omega$ is the area of the surface, and $\delta_{z,z_{k}}\equiv h^{-1}\delta(z-z_{k})$.
The Hamiltonian of point vortices reads
$H=-\frac{\rho_{s}}{2}\sum_{i\neq j}\kappa_{i}\kappa_{j}G(z_{i},z_{j})-\frac{\rho_{s}}{2}\sum_{i}\kappa^{2}_{i}R_{\rm robin}(z_{i}),$ (4)
where $\rho_{s}$ is the superfluid (mass) density,
$R_{\rm robin}(z)=\lim_{w\rightarrow z}\left[G(z,w)+\frac{1}{2\pi}\log d(z,w)\right]$ (5)
is the celebrated Robin function and $d(z,w)$ is the geodesic distance between points $w$ and $z$. When $w\rightarrow z$, $d(z,w)=ds+o(|w-z|)=h|w-z|+o(|w-z|)$ and only the leading order term contributes the Robin function.
The kinetic energy of the fluid is
[…]
where $H_{\text{singular}}=(\rho_{s}/4\pi)\sum_{i}\kappa^{2}_{i}\log d(z_{i},z_{i})$. For the point vortex model, $H_{\text{singular}}=\infty$. In practice, this term is regularized by the vortex core size $\xi$ which serves the ultraviolet cut-off of the PVM, and $H_{\text{singular}}=(\rho_{s}/4\pi)\sum_{i}\kappa^{2}_{i}\log\xi$. Hereafter, we set $\rho_{s}=\kappa=1$ for convenience.
III Statistical mechanics of point vortices on a closed surface
On a closed surface, the total vorticity must vanish, namely, $\sum_{i}\kappa_{i}=0$. We consider a system consisting of $N_{+}$ positive vortices and $N_{-}=N_{+}$ negative vortices. The total vortex number is $N=N_{+}+N_{-}$. The model Eq. (4) exhibits Onsager clustered states at high energies in a bounded domain. In clustered phases, the energy $H\sim N^{2}$, hence in order to describe clustered phases in the proper thermodynamics limit, we need to make the rescaling $\sigma_{i}\rightarrow\sigma_{i}/N_{\pm}$ and define the vortex number density and vortex charge density (vorticity) on a curved surface as following:
$n_{\pm}(x^{\mu})$ $=\frac{1}{N_{\pm}\sqrt{\det g_{\mu\nu}}}\sum^{N_{\pm}}_{i}\delta(x^{\mu}-x^{\mu}_{i}),$ (7)
$\sigma(x^{\mu})$ $=\frac{1}{N_{\pm}\sqrt{\det g_{\mu\nu}}}\sum_{i}\sigma_{i}\delta(x^{\mu}-x^{\mu}_{i}).$ (8)
Here $n_{\pm}$ satisfy the normalization condition $\int{\rm d}\Omega\,n_{\pm}=1$ and ${\sigma}(x^{\mu})=n_{+}(x^{\mu})-n_{-}(x^{\mu})$. In the following we consider vortex distributions at large scales and treat $\sigma$, $n_{\pm}$ and $\psi$ as smooth functions at scales much larger than the mean seperation $\ell$.
For given vortex distributions $n_{\pm}$, the von Neumann entropy reads Joyce and Montgomery (1973)
$S=-\int{\rm d}\Omega\left(n_{+}\log n_{+}+n_{-}\log n_{-}\right),$ (9)
and the energy
$E=\frac{1}{2}\int{\rm d}\Omega\ |\nabla\psi(\textbf{r})|^{2}=\frac{1}{2}\int{\rm d}\Omega\ \psi(\textbf{r}){\sigma}(\textbf{r}),$ (10)
where we have used the relation
$\nabla^{2}\psi(\mathbf{r})$ $=-{\sigma}(\mathbf{r}).$ (11)
The most probable vortex density distribution is given by maximizing the entropy Eq. (9) subject to given energy $E$ and vortex number $N_{\pm}$:
$\delta S-\beta\delta E-\mu_{+}\delta N_{+}/N_{+}-\mu_{-}\delta N_{-}/N_{-}=0,$ (12)
where $\beta,\mu_{\pm}$ are Lagrange multipliers and have the interpretation of inverse temperature and chemical potentials in microcanonical ensemble, respectively. The variation equation (12) gives rise
$n_{\pm}$ $=\exp[\mp\beta\psi(r)+\gamma_{\pm}],$ (13)
[…]
$\nabla^{2}\psi(\mathbf{r})=\exp[\beta\psi(r)+\gamma_{-}]-\exp[-\beta\psi(r)+\gamma_{+}].$ (14)
IV Onsager vortex clusters on a sphere
In this work, we focus on Onsager vortex clusters on a sphere.
IV.1 Stereographic coordinates
The Riemannian metric on a sphere is
$ds^{2}=R^{2}d\theta^{2}+R^{2}\sin^{2}\theta d\phi^{2},$ (15)
where $R$ is the radius, $\theta$ is the polar angle and $\phi$ is the azimuthal angle. In spherical coordinates, the fluid velocity $\textbf{u}=\nabla\psi\times\textbf{e}_{r}$, where $\textbf{e}_{r}$ is the radial unit vector.
@@ IV.1 Stereographic coordinates
The projection from the north pole to the $\zeta=0$ plane induces stereographic coordinates $z=x^{1}+ix^{2}$, where $x^{1}$ and $x^{2}$ axes are identified as the $\xi$ and $\eta$ axes, respectively (Fig. 1). For a sphere, stereographic coordinates are isothermal coordinates and are related to the spherical coordinates by
$z=R\cot\left(\frac{\theta}{2}\right)e^{i\phi}.$ (17)
In terms of stereographic coordinates, the metric reads
$h=\frac{4R^{4}}{(R^{2}+|z|^{2})^{2}}.$ (18)
It is sometimes convenient to introduce polar coordinates $z=\rho e^{i\phi}$, where $\rho=\cot(\theta/2)$ and the polar angle $\phi$ is chosen to coincide with the azimuthal angle. The relations between the three coordinate systems are
@@
In stereographic coordinates, the Green’s function reads Kimura (1999)
$G(z,z_{i})$ $=-\frac{1}{2\pi}\log\frac{|z-z_{i}|}{\sqrt{|z-z_{i}|^{2}+|z\bar{z}_{i}+R^{2}|}}$ (22)
and $R_{\rm robin}=(1/2\pi)\log 2$. In spherical coordinates,
$G(\theta,\phi;\theta_{i},\phi_{i})$ $=-\frac{1}{2\pi}\log\sqrt{2R^{2}-2\mathbf{r}\cdot\mathbf{r}_{i}}$ (23)
$=-\frac{1}{2\pi}\log\left(\sqrt{2}R\sqrt{1-\cos\Theta}\right)$
$=-\frac{1}{2\pi}\log\left(2R\sin\frac{\Theta}{2}\right)$
@@
where $\Theta$ is the angle between $\mathbf{r}$ and $\mathbf{r}_{i}$, and $\cos\Theta=\cos\theta\cos\theta_{i}+\sin\theta\sin\theta_{i}\cos(\phi-\phi_{i})$. It is interesting to notice that $\sqrt{2R^{2}-2\mathbf{r}\cdot\mathbf{r}_{i}}$ is the Euclidean distance, but not the geodesic distance, between two points on the sphere.
IV.2 Onset of clustering
The onset of clustering involves global eigen-modes of the Laplace-Bertrami operator. It is then convenient to use spherical coordinates. Hereafter we set $R=1$ for convenience. On a sphere, besides the energy and the number of vortices, the angular momentum $\mathbf{L}=\int\mathbf{r}\times\mathbf{u}$ is also a conserved quantity. In terms of collective variables of vortices, the components of the angular momentum read
$L_{\xi}$ $=\int{\rm d}\Omega\,\sigma\sin\theta\cos\phi,$ (24)
$L_{\eta}$ $=\int{\rm d}\Omega\,\sigma\sin\theta\sin\phi,$ (25)
$L_{\zeta}$ $=\int{\rm d}\Omega\,\sigma\cos\theta.$ (26)

### v2

@@ Abstract
We study Onsager vortex clustered states in a shell-shaped superfluid containing a large number of quantum vortices. In the incompressible limit and at low temperatures, the relevant problem can be boiled down to the statistical mechanics of neutral point vortices confined on a sphere. We analyze rotation free vortex clustered states within the mean field theory in the microcanonical ensemble. We find that the sandwich state, which involves the separating of vortices with opposite circulation and the clustering of vortices with the same circulation around the poles and the equator, is the maximum entropy vortex distribution, subject to zero angular momentum constraint. The dipole momentum vanishes for the sandwich state and the quadrupole tensor serves as an order parameter to characterize the vortex cluster structure. For given finite angular momentum, the equilibrium vortex distribution forms a dipole structure, i.e., vortices with opposite sign are separated and are accumulated around the south and north pole, respectively. The conditions for the onset of clustering, and the exponents associated with the quadrupole moment and the dipole moment as functions of energy, are obtained within the mean field theory. At large energies, we obtain asymptotically exact vortex density distributions using the stereographic projection method, giving rise to the parameter bounds for the vortex clustered states. The analytical predictions are in excellent agreement with microcanonical Monte Carlo simulations.
@@ I Introduction
Previous studies on Onsager vortices mainly focus on systems confined on a flat region. Recent experimental realizations of BECs Aveline et al. (2020) and ultracold atomic bubbles Carollo et al. (2022) in International Space Station open a possibility to investigate a bubble trapped superfluid and vortex physics on curved surfaces experimentally. Curvature and topology of a surface have non-trivial effects on BECs Tononi and Salasnich (2019); Tononi et al. (2020); Tononi and Salasnich (2023), vortex dynamics Bogomolov (1978); CROWDY (2004); Hally (1980); Polvani and Dritschel (1993); Kimura (1999); Newton (2001); Alobaidi et al. (2006); Aref (2007); Vitelli and Turner (2004); Turner et al. (2010); Dritschel and Boatto (2015); Dritschel et al. (2015); Doshi and Gromov (2021); Bereta et al. (2021); Caracanhas et al. (2022); Padavić et al. (2020); Saito and Hayashi (2023); Li and Efimkin (2023); Xiong and Yu (2024), and the Berezinskii-Kosterlitz-Thouless (BKT) transition Tononi et al. (2022). The formation and structure of Onsager vortices also strongly depend on the surface geometry. On a sphere, the distributions of single species of vortices in the background of uniformly distributed vortices with the opposite circulation have been studied in details Kiessling and Wang (2012). Recent numerical simulations of the point vortex model with removing low energy vortex-antivortex pairs show that, different from the giant dipole configuration for vortices confined to a disc Simula et al. (2014); Yu et al. (2016), in a spherical shell trap uniformly distributed neutral vortices form a quadruple cluster Kanai and Guo (2021). However, it is not clear whether the quadruple clustered state is the statistically most favored state for given energy, angular momentum and vortex number. To answer this question requires the study of distributions of two species of vortices on a sphere simultaneously, which has not been well-addressed.
@@
In this paper, we systematically investigate rotation-free clustered states of a neutral vortex system confined on a sphere. We extend the mean field theory of vortex statistical mechanics developed by Joyce and Montgomery Joyce and Montgomery (1973) to a sphere and treat the distributions of the two species of vortices simultaneously. In particular, we analyze the self-consistent equation of the stream function on a sphere near the uniform state and in the high energy limit. We find that: under the constraint of zero angular momentum, the sandwich state where vortices with opposite sign are distributed around the poles and the equator respectively, is the maximum entropy state and has higher statistical weight than the quadrupole state; for finite angular momentum, a giant dipole structure is statistically favored. The sandwich state and the dipole state are characterized by the non-vanishing quadrupole moment and the dipole moment, respectively. The energy dependence and the upper bounds of the quadrupole moment and the dipole moment at low energies and in the high energy limit are also obtained within the mean-field theory. To test the mean-field predictions, we perform the microcanonical Monte Carlo (MC) sampling of finite vortex numbers and find good agreement on the energy dependence of the quadrupole moment and the dipole moment for a system containing a total of $N=1000$ vortices and the values of the moments showing a saturation to the predicted upper bounds.
II Point vortices on a sphere
@@
For a superfluid on a plane, the circulation of a quanutm vortex is quantized in units of circulation quantum $\kappa\equiv 2\pi\hbar/m$ Pitaevskii and Stringari (2016), and the vorticity has a singularity at the vortex core $\bm{r}_{i}$: $\omega(\bm{r})=\nabla\times\bm{u}=\kappa_{i}\delta(\bm{r}-\bm{r}_{i})=\kappa\sigma_{i}\delta(\bm{r}-\bm{r}_{i})$ with sign $\sigma_{i}=\pm 1$ for singly charged vortices. Here $m$ is the atomic mass and $\bm{u}$ is the superfluid velocity. The quantization ensures that the vorticity of a quantum vortex concentrates around the core region Eyink and Sreenivasan (2006). Hence when the mean separation between quantum vortices $\bar{d}$ is much larger than the vortex core size $\xi$ Fetter (1998); Barenghi and Parker (2016), the point vortex model (PVM) describes well the dynamics of quantum vortices in the incompressible limit Onsager (1949); Eyink and Sreenivasan (2006); Novikov (1975); Aref et al. (1999), provided vortex annihilation can be neglected. In this limit, the flow is nearly incompressible $\nabla\cdot\bm{u}=0$ and hence a stream function $\psi$ can be introduced such that $\bm{u}=\nabla\psi\times\hat{\bm{z}}$, where $\hat{\bm{z}}$ is the unit vector normal to the plane. The stream function is completely determined by vortex positions and $\psi(\bm{r})=-\sum_{i}\kappa_{i}G(\bm{r},\bm{r}_{i})$, where $G(\bm{r},\bm{r}_{i})=-(1/2\pi)\log(|\bm{r}-\bm{r}_{i}|/\xi)$ is the Green’s function of Laplacian operator in 2D: $\nabla^{2}G(\bm{r},\bm{r}_{i})=-\delta(\bm{r}-\bm{r}_{i})$. The PVM also describes 2D superfluid transition at finite temperature, which is the celebrated BKT transition Kosterlitz and Thouless (1973); Kosterlitz (1974).
@@
On curved surfaces, a covariant PVM has been well formulated Hally (1980); Kimura (1999); Dritschel and Boatto (2015). Here we only focus on the PVM on a sphere. The Riemannian metric on a sphere is
${\rm d}s^{2}=R^{2}{\rm d}\theta^{2}+R^{2}\sin^{2}\theta{\rm d}\phi^{2},$ (1)
where $R$ is the radius, $\theta$ is the polar angle and $\phi$ is the azimuthal angle.
@@
The projection from the north pole to the $\zeta=0$ plane induces stereographic coordinates $z=x^{1}+{\rm i}x^{2}$, where $x^{1}$ and $x^{2}$ axes are identified as the $\xi$ and $\eta$ axes, respectively (Fig. 1). Let us also introduce the following notations $\partial_{z}\equiv(\partial_{1}-{\rm i}\partial_{2})/2$ and $\partial_{\bar{z}}\equiv(\partial_{1}+{\rm i}\partial_{2})/2$ . In terms of stereographic coordinates, the metric reads ${{\rm d}}s^{2}=h(x^{1},x^{2})[({\rm d}x^{1})^{2}+({\rm d}x^{2})^{2}]$, where $h=4R^{4}/(R^{2}+|z|^{2})^{2}$. For a sphere, stereographic coordinates are isothermal coordinates Chern (1955) and are related to the spherical coordinates by
$z=R\cot\left(\frac{\theta}{2}\right){\rm e}^{\rm i\phi}.$ (3)
It is sometimes convenient to introduce polar coordinates $z=\rho{\rm e}^{\rm i\phi}$, where $\rho=R\cot(\theta/2)$ and the polar angle $\phi$ is chosen to coincide with the azimuthal angle. The relations between the three coordinate systems are
@@
In spherical coordinates, the stream function reads
$\psi(\theta,\phi)=\kappa\sum_{i}\sigma_{i}G(\theta,\phi;\theta_{i},\phi_{i}),$ (7)
where $(\theta_{i},\phi_{i})$ is the position of the $i$th vortex, $G(\theta,\phi;\theta_{i},\phi_{i})$ is the Green’s function satisfying Kimura (1999); Dritschel and Boatto (2015)
$\nabla^{2}G(\theta,\phi;\theta_{i},\phi_{i})=-\delta_{s}(\theta,\phi;\theta_{i},\phi_{i})+1/\Omega,$ (8)
where
$\delta_{s}(\theta,\phi;\theta_{i},\phi_{i})=\frac{1}{R^{2}\sin\theta}\delta(\theta-\theta_{i})\delta(\phi-\phi_{i}),$ (9)
$\Omega=4\pi R^{2}$ is the area of the sphere, and $\nabla^{2}=(1/\sin\theta)\partial_{\theta}(\sin\theta\partial_{\theta})+(1/\sin^{2}\theta)\partial^{2}_{\phi}$ is the Laplace-Bertrami operator. The Green’s function reads Kimura (1999)
$G(\theta,\phi;\theta_{i},\phi_{i})$ $=-\frac{1}{2\pi}\log\frac{\sqrt{2R^{2}-2\bm{r}\cdot\bm{r}_{i}}}{\xi}$ (10)
$=-\frac{1}{2\pi}\log\left(\frac{\sqrt{2}R}{\xi}\sqrt{1-\cos\Theta}\right),$
@@
where $\Theta$ is the angle between $\bm{r}$ and $\bm{r}_{i}$, and $\cos\Theta=\cos\theta\cos\theta_{i}+\sin\theta\sin\theta_{i}\cos(\phi-\phi_{i})$. It is interesting to notice that $\sqrt{2R^{2}-2\bm{r}\cdot\bm{r}_{i}}$ is the Euclidean distance, but not the geodesic distance, between two points on the sphere.
On a closed surface, the total vorticity must vanish, namely, $\sum_{i}\sigma_{i}=0$. Hence combining Eqs. (7) and (8), we obtain
$\nabla^{2}\psi(\theta,\phi)=-\omega(\theta,\phi),$ (11)
where
$\omega(\theta,\phi)=\kappa\sum_{i}\sigma_{i}\delta(\theta,\phi;\theta_{i},\phi_{i}).$ (12)
The Hamiltonian of the point vortices on a sphere is
$H=\frac{\rho_{s}\kappa^{2}}{2}\sum_{i\neq j}\sigma_{i}\sigma_{j}G(\theta_{i},\phi_{i};\theta_{j},\phi_{j}).$ (13)
On a sphere, besides the energy and the number of vortices, the following quantities
$L_{\xi}$ $=\kappa R^{2}\sum_{i}\sigma_{i}\sin\theta_{i}\cos\phi_{i},$ (14)
$L_{\eta}$ $=\kappa R^{2}\sum_{i}\sigma_{i}\sin\theta_{i}\sin\phi_{i},$ (15)
$L_{\zeta}$ $=\kappa R^{2}\sum_{i}\sigma_{i}\cos\theta_{i},$ (16)
are also conserved and are associated with the fluid angular momentum.
III Onsager vortex clusters on a sphere
III.1 Onsager vortex clusters at negative temperature
A particular feature of 2D point-vortex systems in a bounded domain is that the phase space area per vortex is finite. For a vortex system on a sphere, the phase space of a vortex is the spherical surface. Let us subdivide the surface into small cells of area $a_{i}\ll\Omega$, where $\Omega$ is the surface area. While the cells are still large enough to contain many vortices. We consider a system consisting of $N_{+}$ positive vortices and $N_{-}=N_{+}$ negative vortices. The total vortex number is $N=N_{+}+N_{-}$. We denote the number of vortices in a cell of each species as $N_{i,\hat{\sigma}}$ with $\hat{\sigma}$$\pm$ and the number of states is then given by Joyce and Montgomery (1973); Maestrini and Salman (2019)
[…]
Then the thermodynamic entropy is $k_{B}\log\mathcal{W}(H)$ , where $k_{B}$ is the Boltzmann constant. Since $k_{B}\log\mathcal{W}(-\infty)=k_{B}\log\mathcal{W}(+\infty)=0$, which correspond to extreme situations of vortex-antivortex pairs collapsing and like-sign vortices concentrating into a point, respectively. Then $\mathcal{W}$ must reach its maximum at an intermediate energy $H_{m}$ and the inverse temperature $1/T\equiv k_{B}\partial\log\mathcal{W}(H)/\partial H$ is thus negative for $H>H_{m}$. High energy states at negative temperature correspond to the formation of macroscopic vortex clusters Onsager (1949); Eyink and Sreenivasan (2006).
For given total vortex numbers $N_{\pm}$ and the area $\Omega$, the entropy is a function of $\{N_{i,\pm}\}$ and for large $N_{i,\pm}$ we obtain
$(1/N_{\pm})\log\mathcal{W}$ $\simeq-\sum_{i}\left(\frac{N_{i,+}}{N_{+}a_{i}}\log\frac{N_{i,+}}{N_{+}a_{i}}\right)a_{i}$ (18)
$-\sum_{i}\left(\frac{N_{i,-}}{N_{-}a_{i}}\log\frac{N_{i,-}}{N_{-}a_{i}}\right)a_{i},$
where the Stirling’s formula is applied and the irrelevant constants are ignored.
III.2 Continuous limit and the mean-field theory
The PVM Eq. (13) exhibits Onsager clustered states at high energies. In clustered phases, the energy $H\sim N^{2}$, hence in order to describe clustered phases in the proper thermodynamics limit, we need to make the rescaling $\sigma_{i}\rightarrow\sigma_{i}/N_{\pm}$ and define the vortex number densities on a sphere as following:
$n_{\pm}(\bm{r})$ $=\frac{1}{N_{\pm}}\sum^{N_{\pm}}_{i}\delta_{s}(\theta,\phi;\theta_{i},\phi_{i}),$ (19)
where $n_{\pm}$ satisfy the normalization condition $\int{\rm d}\Omega\,n_{\pm}=1$. Then the rescaled vortex charge density (vorticity) is
$\omega(\bm{r})=\kappa{\sigma}(\bm{r})\equiv\kappa[n_{+}(\bm{r})-n_{-}(\bm{r})],$ (20)
and the rescaled stream function
$\psi(\bm{r})=\kappa\int{\rm d}\Omega^{\prime}G(\bm{r},\bm{r}^{\prime})\,\sigma(\bm{r}^{\prime}),$ (21)
satisfying
$\nabla^{2}\psi(\bm{r})=-\kappa{\sigma}(\bm{r}).$ (22)
In the following we consider vortex distributions at large scales and treat $\sigma$, $n_{\pm}$ and $\psi$ as smooth functions at scales much larger than the mean seperation $\bar{d}$. The rescaled coarse-grained Hamiltonian reads
$E$ $=\frac{\rho_{s}\kappa^{2}}{2}\int{\rm d}\Omega{\rm d}\Omega^{\prime}\sigma(\bm{r})G(\bm{r},\bm{r}^{\prime})\sigma(\bm{r}^{\prime})$ (23)
$=\frac{\rho_{s}}{2}\int{\rm d}\Omega\,\omega(\bm{r})\psi(\bm{r}),$
and $E\sim\mathcal{O}(1)$. For a sphere, this Hamiltonian is nothing but the rescaled kinetic energy of the flow generated by vortices: $E=\rho_{s}/2\int{\rm d}\Omega\,|\bm{u}|^{2}=\rho_{s}/2\int{\rm d}\Omega\,|\nabla\psi|^{2}=\rho_{s}/2\int{\rm d}\Omega\,\omega\,\psi$, where the fluid velocity $\bm{u}=\nabla\psi\times\bm{e}_{r}$ and $\bm{e}_{r}$ is the radial unit vector. Applying point vortices directly to Eq. (23) [plugging Eqs. (7) and (12) into Eq. (23)], we obtain that $E=H/N^{2}_{\pm}+H_{\text{singular}}$, where $H_{\text{singular}}=(\rho_{s}\kappa^{2}/4\pi N^{2}_{\pm})\sum_{i}\sigma^{2}_{i}\log[d(\bm{r}_{i},\bm{r}_{i})/\xi]$, and $d(\bm{r}_{i},\bm{r}_{j})$ is the Euclidean distance between the two points on a sphere. For point vortices, $H_{\text{singular}}=\infty$. In practice, the distance between vortices can not be smaller than the vortex core size $\xi$ which serves the ultraviolet cut-off of the PVM. By choosing $d(\bm{r}_{i},\bm{r}_{i})=\xi$, we have $H_{\text{singular}}=0$.
In terms of the rescaled collective variables of vortices, the components of the angular momentum read
$L_{\xi}$ $=\kappa R^{2}\int{\rm d}\Omega\,\sigma\sin\theta\cos\phi,$ (24)
$L_{\eta}$ $=\kappa R^{2}\int{\rm d}\Omega\,\sigma\sin\theta\sin\phi,$ (25)
$L_{\zeta}$ $=\kappa R^{2}\int{\rm d}\Omega\,\sigma\cos\theta.$ (26)